Solve and check for X :
[5/(x+6)]+[2/(x^2+7x+6)]+[3/(x+1)]

Answer by orca(409)

Do you mean solving the following equation for x?
[5/(x+6)]+[2/(x^2+7x+6)]+[3/(x+1)] = 0
If so, see the solution below:
SOLUTION:
note that x^2+7x+6 = (x+1)(x+6)
[5/(x+6)]+[2/(x+1)(x+6)]+[3/(x+1)]= 0
Multiplying both sides by (x+1)(x+6), we have
5(x+1) + 2 +3(x+6) = 0
5x + 5 + 2 + 3x + 18 = 0
8x + 25 = 0
x = -25/8